What is the advantage of knowing that two ? Calc
Answer #1

2. Fe3+ (aq) + 3 e = Fe (s)

E0Fe3+/Fe = - 0.04 V

3. [Cr2O7]2-

2 x + 7( -2 ) = - 2

2x - 14 = -2  

2x = -2 +14.

2x = 12.

x = oxiation stateof Cr = +6

and oxidation state of Cr in Cr3+ = +3.

4.

Fe(s) / Fe3+ (aq) || MnO42-(aq)/ Mn2+(aq)